OpenStudy (anonymous):

Express log_4 27 - 2(log_4 3) as a single logarithm.

OpenStudy (lucaz):

you know the log properties?

OpenStudy (anonymous):

I don't really understand them

OpenStudy (anonymous):

I'm thinking it is either log_4 12 or log_4 18

OpenStudy (lucaz):

I'm gonna attach a file

OpenStudy (anonymous):

okay!

OpenStudy (mathmale):

Ricardo: Do you happen to have a reference of any kind (book, Internet or whatever) that lists the three main properties of logs that apply to this kind of question? If so, that'd be really helpful.

OpenStudy (lucaz):

yes, they are not difficult to use, just take a lookon google

OpenStudy (anonymous):

I think I understand it a bit more. I am terrible at this sort of thing.

OpenStudy (lucaz):

it's the same as the distributive property, a(b+c) = ab + ac, you just have to follow the rules

OpenStudy (anonymous):

Yeah! I see that now! Can you help me with this one? Solve 3^(2x) = 30. log_3 (30/2) log_3 15 2(log_3 30) log_3 60

OpenStudy (mathmale):

R: take the log of both sides. You certainly can, if you wish, take the log to base 3 of both sides, ending up with 2x=log_3 30. How muich sense does this make to you?

OpenStudy (anonymous):

It makes a good amount of sense. So now if I have 2x=log_3 30 I would divide boths sides by 2x and get x = log_3 15 ?

OpenStudy (mathmale):

Yes, divide both sides by 2; but the right side ends up as (1/2)log_3 30. Can't divide that 30 by 2. Make a note of this so we can discuss it in more depth when we again have access to Equation Editor and Draw.

OpenStudy (anonymous):

Okay, but that answer isn't one of my answer choices that they had given me..

OpenStudy (mathmale):

If that's the case, I'd bet your answer choices involve either ln or log, not log_3. Correct?

OpenStudy (anonymous):

a. log_3 (30/2) b. log_3 15 c. 2(log_3 30) d. log_3 60 Those are the answer choices.

OpenStudy (lucaz):

it should be C but 1/2( ) not 2( )

OpenStudy (mathmale):

Solve 3^(2x) = 30. Take log_3 of both sides, resulting in 2x=log_3 30 Solve for x: [log_3 30]/2 = x. I agree with lucaz.

OpenStudy (anonymous):

That's what I was thinking. Maybe the quiz had a glitch or they forgot to put that in.

OpenStudy (mathmale):

Ric: Again, I 'd suggest you and I (or you and others) practice application of these and other rules of logs once we again have access to Equation Editor and Draw. Care to discuss any other practice problems right now?

OpenStudy (anonymous):

Nope, All of the rest I know I can do on my own. Thank you to the both of you!

OpenStudy (lucaz):

no problem..

OpenStudy (mathmale):

You're welcome!